The bottom of a 4.0-m-tall diving bell is at an unknown depth underwater. The pressure of the air inside the bell is 2.0 atm (it was 1 atm before the bell entered the water). The average density of ocean water is slightly greater than the density of fresh water, ρocean water = 1027 kg/m 3. How high is the water inside the bell and how deep is the bottom of the bell under the water?


1
Expert's answer
2021-05-31T15:38:24-0400

a)


"h=\frac{1}{2}4=2\ m"

b)


"P'=P+\rho g(H-h)\\2(101325)=101325+(1027)(9.8)(H-2)\\H=12.1\ m"